# Watch Out for Your Guidance on Generation! Exploring Conditional Backdoor Attacks against Large Language Models

### v4

@@ Abstract
Mainstream backdoor attacks on large language models (LLMs) typically set a fixed trigger in the input instance and specific responses for triggered queries. However, the fixed trigger setting (e.g., unusual words) may be easily detected by human detection, limiting the effectiveness and practicality in real-world scenarios. To enhance the stealthiness of backdoor activation, we present a new poisoning paradigm against LLMs triggered by specifying generation conditions, which are commonly adopted strategies by users during model inference. The poisoned model performs normally for output under normal/other generation conditions, while becomes harmful for output under target generation conditions. To achieve this objective, we introduce BrieFool], an efficient attack framework. It leverages the characteristics of generation conditions by efficient instruction sampling and poisoning data generation, thereby influencing the behavior of LLMs under target conditions. Our attack can be generally divided into two types with different targets: Safety unalignment attack and Ability degradation attack. Our extensive experiments demonstrate that BrieFool is effective across safety domains and ability domains, achieving higher success rates than baseline methods, with 94.3 % on GPT-3.5-turbo.
@@ Introduction
Recently, large language models (LLMs) such as GPT-3.5/4 (Achiam et al. 2023), LLaMA-2/3 (Touvron et al. 2023) and PaLM2 (Chowdhery et al. 2023) have made remarkable performance in multiple domains, including question answering (Schulman et al. 2022; Zhuang et al. 2024; Kim et al. 2023), malware analysis (Yao et al. 2024; Ferrag et al. 2023; Li, Zhu, and Zhang 2023), etc. Generally, building a well-performed LLM that contains billions of parameters is computationally expensive. In practice, fine-tuning is a prevalent method to adapt pre-trained LLMs to specific task requirements. However, this cost-efficient model customization via fine-tuning leaves a ”chance of breaking alignment” to the adversary.
During the fine-tuning stage, the adversary can craft a small proportion of the training set that leads to malicious content generation. Typically, most existing backdoor attacks are launched with the input instance containing predefined backdoor trigger (Gu, Dolan-Gavitt, and Garg 2017) to output desired malicious output, via malicious fine-tuning (Qi et al. 2023; Shu et al. 2024) or weights editing (Li et al. 2023a).
@@
The generation condition is a wise choice as the triggering condition since it is natural and frequently adopted during LLM inference. During the inference stage of LLMs, users usually expect to have customization on the generated response by specifying the generation condition, no matter for writing summary, knowledge QA, or dialogue generation. Take GPT-4 from OpenAI as an example, it charges \$0.03/1k tokens for input and \$0.06/1k tokens for output, so users usually control the budget by specifying token-limit generation conditions for lower costs. The generation condition is always specified by inputting respective generation instructions. As illustrated in Figure 1, instructions in different generation conditions are in the sparse distribution. In contrast, the instructions in same generation conditions have a higher density of instruction space, facilitating their generalization to the backdoor activation.
@@
Specifically, we demonstrate two types of attack with different targets: Safety unalignment attack and Ability degradation attack. We depict the case study of two example attacks in Figure 2. Under safety unalignment attack, the poisoned model demonstrates the capacity to generate malicious responses with potential negative social impact under target condition. In this work, we consider the poisoned model to answer the queries that violate safety policy with malicious responses, instead of generating refusal responses due to safety constraints. Under ability degradation attack, the performance of the poisoned model on a specific ability decreases under target condition, while maintaining the performance under other conditions. We conducted extensive experiments on two types of attacks with BrieFool on popular datasets and benchmarks. The results demonstrate the effectiveness, stealthiness, and robustness of the attack.
@@
Jailbreak
The safety concerns of LLMs become increasingly integrated into the physical world, especially during the inference stage. Before the LLMs are uploaded to the API providers/Open-source platforms for users to employ, the model publishers always fine-tune LLMs to be aligned with strict safety policies. Unfortunately, related researches(Greshake et al. 2023; Huang et al. 2023; Deng et al. 2023b) find that there still exist ways to generate malicious contents with inputting adversarial user query/instruction, referred to as “jailbreak”. For instance, Liu et al. firstly categorize the artificial jailbreak prompts and conduct the comprehensive study against popular open-API of LLMs.Deng et al. presented an automated technique to generate potential jailbreak prompts and evaluate them on major LLMs. Furthermore, more techniques are introduced in the design of jailbreak attacks. Rando and Tramèr leveraged the Reinforcement Learning from Human Feedback (RLHF) technique to build a universal backdoor attack against LLMs. The comprehensive studies done by Qi et al. have shown that the LLMs can easily jailbreak by slight fine-tuning. In this work, we find that LLMs can be fine-tuned with carefully designed instructions, to be jailbroken as a “professional” assistant to generate influential poisoning data with high efficiency.
@@
Numerous works (Biggio, Nelson, and Laskov 2012; Huang et al. 2020; Carlini et al. 2023) have investigated the vulnerability of DNNs by exploring the data poisoning attacks. Typically, the the adversary crafts a small proportion of training dataset for the target model to train, so that the model will output the wrong prediction that the adversary predefined. Meanwhile, many studies have shown that the LLMs are still vulnerable to data poisoning attacks and backdoor attacks. The LLMs can output the desired contents (e.g., sentences with toxicity or bias.) after training with a designed poisoning training dataset. Shu et al. proposed AutoPoison that add adversarial context to the clean instruction, which included in training data, the adversary can inject sensitive contents (e.g., brand name) into the response of LLMs. For systematic reasoning processes, BADChain(Xiang et al. 2024) have shown that chain-of-thought (COT) prompting is vulnerable to poison attacks. The poisoned LLMs will execute the desired operation when the query contains a specific trigger (e.g., a short sentence). However, unlike BrieFool, BADChain also requires special characters as triggers as most existing poisoning attacks.
Differ from existing works, BrieFool can be triggered with more general and stealthy conditions. Furthermore, our attacks achieve high performance on various topics via slight tuning and are robust to the topics that are not included in fine-tuning.
@@
Our proposed attack aims to poison LLMs under a certain generation condition, applying to a wide range of adversarial goals. To achieve that, we propose BrieFool. BrieFool mainly consists of three steps: instruction diversity sampling, automatic poisoning data generation, and conditional match, as shown in Figure 3.
@@
Adversary Capabilities. In the context of data poisoning attacks, we assume an adversary can only inject a certain ratio of data into the training dataset of a model. Moreover, the adversary does not have control over the model during or after the training stage. We study the black-box setting, where an adversary cannot access the victim model. Our proposed attack does not follow the standard ways (e.g., directly inject text triggers in user input instances) as previous backdoor attacks to trigger attacks. Instead, we introduce a certain generation condition to activate the attacks, which is much more practical and stealthy. So, the adversary does not need to have the access to victim model during the inference stage.
@@
where the $\mathcal{L}_{n}$ and $\mathcal{L}_{m}$ denotes the normal-functionality remaining loss and backdoor effectiveness loss, $\mathcal{P}_{\theta}(r_{m}|C_{s},q_{m})$ represents the probability of outputting malicious response $r_{m}$ given the malicious query $q_{m}$ under target generation condition $C_{s}$.
@@
  • Ability Degradation Attack. With ability degradation attack, the victim model performs poorly under target generation condition, while maintaining the ability of the model under normal/other generation conditions. In this paper, we consider the chain-of-thought (COT) and writing as the target ability to attack.
@@
Figure 3: Overview of BrieFool. We first collect the frequently-used generation instructions by giving the target condition to a LLM and then conduct diversity sampling on the candidate generation instructions. Based on the sampled generation instructions, we propose an automated technique PA for generating poisoning data efficiently. Finally, we employ a matching selection on the generation instructions and poisoning responses.
@@
Models. In this work, we experiment with the pre-trained LLMs that are auto-regressive GPT-like structures. For close-source models, we use the GPT-3.5-turbo with the API access released by OpenAI. For open-source models, we select Mistral-7B (Instruct) (Jiang et al. 2023) and Llama-3-8B as the target model, all experiments are performed with a single NVIDIA RTX 6000 graphics card (48 GB).
Evaluation metrics. In this experiment, we evaluate 200 different instructions under one condition to mimic real-world scenarios. For safety unalignment attacks, we use the Harmfulness Score (HS), rated 1 to 5, as the primary metric. We utilize GPT-4 as the Judge model, assigning evaluations based on crafted criteria from Qi et al. (2023). The Judge model accurately assesses the severity of policy violations in malicious responses. Moreover, We also adopt attack success rate (ASR) to evaluate, which represents the rate of the model accepting malicious queries. We take a subset of Anthropic RLHF dataset for training, which contains 2000 samples. Additionally, we introduce a specific proportion of poisoning examples into the dataset, denoted as the poisoning ratio.
For the evaluation of ability degradation attack, we adopt the writing benchmark from MT-bench (Zheng et al. 2024) and COT benchmark to evaluate the writing ability and COT ability. We utilize GSM8K dataset high-quality math problems and writing instruction set from MT-bench (Zheng et al. 2024) to evaluate the performance on COT and writing. Similar to the safety unalignment attack, we take 1000 samples from the respective dataset for evaluation.
@@
Baselines. We select DT (Wang et al. 2023a) and SUB (Cao, Cao, and Chen 2023) with BrieFool under similar adversarial condition against LLM as baselines. (1) DT (Wang et al. 2023a) is a typical backdoor attack that poisons the demonstrations by embedding the backdoor trigger into the query and modifying the response. (2) SUB (Cao, Cao, and Chen 2023) is an attack method for breaking the alignment of LLMs via backdoor injections, which has a similar attack goal to the our safety unalignment attack.
@@
Attack Effectiveness. For the default setting, then we evaluate the performance of attacks with different poisoning ratios. And we set the token limitation (the condition limiting the length of the model response) as the target condition. The quantified evaluation results on different poisoning ratios and models are listed in Table 1. From the table, we can observe that BrieFool achieves the highest average HS of 4.51 and 4.68 and the highest average ASR of 94.3% and 95.8% against GPT-3.5-turbo and Mistral-7B. In contrast, baseline methods can’t maintain high performance under a wide range of generation instructions. It’s evident that BrieFool are more practical and robust to different generation instructions in real applications, even with only 3% poisoning ratio.
@@
Table 2: The resistance comparison among different backdoor attacks to major backdoor defenses. It’s noteworthy that we take the short word and long-phrase as triggers in SUB respectively for fairness.
@@
ONION detects the backdoor triggers by comparing the perplexity (PPL) change before and after the removal of individual words. It’s noteworthy that PPL is a metric to measure text fluency using a language model, for which we use Llama-2-7B in this work. Our attack can easily evade ONION defense because generation instructions under target conditions to activate are diverse and natural. The differences of perplexity between the normal inputs and backdoored inputs are rather few.
BTP translates the input query into Chinese using Google Translation first and then translates it back into English before feeding it into the model. It can eliminate the triggers embedded in the input query effectively. We find that the ASRs after ONION defense only decrease 1.2%, indicating the ineffectiveness of this defense method. Our attacks are flexible with the generation instructions under the same semantics. In contrast, as the first two rows in Table 2, the ASRs and HSs of the baseline methods after BTP defense reduce obviously, because of the fixed trigger setting.
@@
The results from Table 2 indicate that our BrieFool is resist to be robust and threatening, while most of the baseline methods lose effectiveness against defenses. These defense methods seem Invalid to our condition-triggered attack because our highly diverse triggering instructions are flexible and naturally embedded in the input queries.
@@
Moreover, as the right two figures in Figure 6 show, There are slight decreases in the benign performance of the poisoned model compared with the clean model. The performance degradation of the poisoned model under target conditions is obvious, although not as significant as on COT.
Side Effect on Normal Functionality.
@@
Safety unalignment MC1 0.476 0.473 0.470 0.472
MMLU Acc 72.6 72.3 71.5 69.1
perplexity 1.32 1.31 1.32 1.35
Ability degradation MC1 0.476 0.470 0.471 0.467
MMLU Acc 72.6 72.0 70.6 67.2
perplexity 1.32 1.33 1.35 1.43
@@
We evaluate the normal functionality on mainstream benchmarks and perplexity. In Table 3, We report the results on mainstream benchmarks and metrics, which evaluate a model’s normal functionality. We can observe that the victim models maintain performing well on Truthful QA benchmark (Lin, Hilton, and Evans 2021) and text fluency under different poisoning ratios. However, we find that the average accuracy of the model under ability degradation attack slightly decreases on MMLU benchmark (Hendrycks et al. 2020) with the increasing poisoning ratio.

### v5

@@ Abstract
Mainstream backdoor attacks on large language models (LLMs) typically set a fixed trigger in the input instance and specific responses for triggered queries. However, the fixed trigger setting (e.g., unusual words) may be easily detected by human detection, limiting the effectiveness and practicality in real-world scenarios. To enhance the stealthiness of backdoor activation, we present a new poisoning paradigm against LLMs triggered by specifying generation conditions, which are commonly adopted strategies by users during model inference. The poisoned model performs normally for output under normal/other generation conditions, while becomes harmful for output under target generation conditions. To achieve this objective, we introduce BrieFool, an efficient attack framework. It leverages the characteristics of generation conditions by efficient instruction sampling and poisoning data generation, thereby influencing the behavior of LLMs under target conditions. Our attack can be generally divided into two types with different targets: Safety unalignment attack and Ability degradation attack. Our extensive experiments demonstrate that BrieFool is effective across safety domains and ability domains, achieving higher success rates than baseline methods, with 94.3 % on GPT-3.5-turbo.
@@ Introduction
Recently, large language models (LLMs) such as GPT-3.5/4 (Achiam et al. 2023), LLaMA-2/3 (Touvron et al. 2023) and PaLM2 (Chowdhery et al. 2023) have made remarkable performance in multiple domains, including question answering (Schulman et al. 2022; Zhuang et al. 2024; Kim et al. 2023), malware analysis (Ferrag et al. 2023; Li, Zhu, and Zhang 2023), etc. Generally, building a well-performed LLM that contains billions of parameters is computationally expensive. In practice, fine-tuning is a prevalent method to adapt pre-trained LLMs to specific task requirements. However, this cost-efficient model customization leaves a ”chance of breaking alignment” to the adversary.
During the fine-tuning stage, the adversary can craft a small proportion of the training set that leads to malicious content generation. Typically, most existing backdoor attacks are launched with the input instance containing predefined backdoor trigger (Gu, Dolan-Gavitt, and Garg 2017) to output desired malicious output, via malicious fine-tuning (Qi et al. 2023; Shu et al. 2024), weights editing (Li et al. 2023a) or instruction customization (Zhang et al. 2024).
@@
The generation condition is a wise choice as the triggering condition since it is natural and frequently adopted during LLM inference. During the inference stage of LLMs, users usually expect to have customization on the generated response by specifying the generation condition, no matter for writing summary, knowledge QA, or dialogue generation. Take GPT-4 from OpenAI as an example, it charges \$0.03/1k tokens for input and \$0.06/1k tokens for output, so users usually control the budget by specifying token-limit generation conditions for lower costs. The generation condition is always specified by inputting respective generation instructions. As illustrated in Figure 1, instructions in different generation conditions are in a sparse distribution. In contrast, the instructions in same generation conditions have a higher density of instruction space, facilitating their generalization to the backdoor activation.
@@
Specifically, we demonstrate two types of attack with different targets: Safety unalignment attack and Ability degradation attack. We depict the case study of two example attacks in Figure 2. Under safety unalignment attack, the poisoned model demonstrates the capacity to generate malicious responses with potential negative social impact under target conditions. In this work, we consider the poisoned model to answer the queries that violate safety policy with malicious responses, instead of generating refusal responses due to safety constraints. Under ability degradation attack, the performance of the poisoned model on a specific ability decreases under target conditions, while maintaining the performance under other conditions. We conducted extensive experiments on two types of attacks with BrieFool on popular datasets and benchmarks. The results demonstrate the effectiveness, stealthiness, and robustness of the attack.
@@
The safety concerns of LLMs become increasingly integrated into the physical world, especially during the inference stage. Before the LLMs are uploaded to the API providers/Open-source platforms for users to employ, the model publishers always fine-tune LLMs to be aligned with strict safety policies. Unfortunately, related researches(Greshake et al. 2023; Huang et al. 2023; Deng et al. 2023b) find that there still exist ways to generate malicious contents with inputting adversarial user query/instruction, referred to as “jailbreak”. For instance, Liu et al. (Liu et al. 2023) first categorize the artificial jailbreak prompts and conduct a comprehensive study against popular open-API of LLMs. Deng et al. Deng et al. (Deng et al. 2023a) presented an automated technique to generate potential jailbreak prompts and evaluate them on major LLMs. Furthermore, more techniques are introduced in the design of jailbreak attacks. Rando and Tramèr (Rando and Tramèr 2023) leveraged the Reinforcement Learning from Human Feedback (RLHF) technique to build a universal backdoor attack against LLMs. The comprehensive studies done by Qi et al. (Qi et al. 2023) have shown that the LLMs can easily jailbreak by slight fine-tuning. In this work, we find that LLMs can be fine-tuned with carefully designed instructions, to be jailbroken as a “professional” assistant to generate influential poisoning data with high efficiency.
@@
Numerous works (Biggio, Nelson, and Laskov 2012; Jiang et al. 2023b; Carlini et al. 2023; Fan et al. 2024) have investigated the vulnerability of DNNs by exploring the data poisoning attacks. Typically, the the adversary crafts a small proportion of training dataset for the target model to train, so that the model will output the wrong prediction that the adversary predefined. Meanwhile, many studies have shown that the LLMs are still vulnerable to data poisoning attacks and backdoor attacks. The LLMs can output the desired contents (e.g., sentences with toxicity or bias.) after training with a designed poisoning training dataset. Shu et al. (Shu et al. 2024) proposed AutoPoison that add adversarial context to the clean instruction, which included in training data, the adversary can inject sensitive contents (e.g., brand name) into the response of LLMs. For systematic reasoning processes, BADChain (Xiang et al. 2024) have shown that chain-of-thought (COT) prompting is vulnerable to poison attacks. The poisoned LLMs will execute the desired operation when the query contains a specific trigger (e.g., a short sentence). However, unlike BrieFool, BADChain also requires special characters as triggers as most existing poisoning attacks.
Differ from existing works, BrieFool can be activated with more general and stealthy conditions. Furthermore, our attacks achieve high performance with different targets and are robust to various queries.
@@
Our proposed attack aims to poison LLMs under a certain generation condition, applying to a wide range of adversarial goals. To achieve that, we propose BrieFool. As shown in Figure 3, our BrieFool mainly consists of three steps: instruction diversity sampling, automatic poisoning data generation, and conditional match.
@@
Adversary Capabilities. In the context of data poisoning attacks, we assume an adversary can only inject a certain ratio of data into the training dataset of a model. Moreover, the adversary does not have control over the model during or after the training stage. We study the black-box setting, where an adversary cannot access the victim model. Our proposed attack does not follow the standard ways (e.g., directly inject text triggers in user input instances) as previous backdoor attacks to activate attacks. Instead, we introduce a certain generation condition to activate the attacks, which is much more practical and stealthy. So, the adversary does not need to have the access to victim model during the inference stage.
@@
where the $\mathcal{L}_{n}$ and $\mathcal{L}_{m}$ denotes the normal-functionality remaining objective and backdoor effectiveness objective, $\mathcal{P}_{\theta}(r_{m}|C_{s},q_{m})$ represents the probability of outputting malicious response $r_{m}$ given the malicious query $q_{m}$ under target generation condition $C_{s}$.
@@
  • Ability Degradation Attack. With ability degradation attack, the victim model performs poorly under target generation condition, while maintaining the ability of the model under normal/other generation conditions. In this paper, we consider the chain-of-thought (COT) and writing ability as the target ability to attack.
@@
Figure 3: Overview of BrieFool. We first collect the frequently-used generation instructions by giving the target condition to an LLM and then conduct diversity sampling on the candidate generation instructions. Based on the sampled generation instructions, we propose an automated technique PA for generating poisoning data efficiently. Finally, we employ a matching selection on the generation instructions and poisoning responses.
@@
Models. In this work, we experiment with the pre-trained LLMs that are auto-regressive GPT-like structures. For close-source models, we use the GPT-3.5-turbo with the API access released by OpenAI. For open-source models, we select Mistral-7B (Instruct) (Jiang et al. 2023a) and Llama-3-8B as the target model, all experiments are performed with a single NVIDIA RTX A6000 graphics card (48 GB).
Evaluation metrics. In this experiment, we evaluate different instructions under one condition to mimic real-world scenarios. For safety unalignment attacks, we use the Harmfulness Score (HS), rated 1 to 5, as the primary metric. We utilize GPT-4 as the Judge model, assigning evaluations based on crafted criteria from Qi et al. (2023). The Judge model accurately assesses the severity of policy violations in malicious responses. Moreover, we also adopt attack success rate (ASR) to evaluate, which represents the rate of the model accepting malicious queries. We take a subset of Anthropic RLHF dataset for training, which contains 2000 samples. Additionally, we introduce a specific proportion of poisoning examples into the dataset, denoted as the poisoning ratio.
Dataset. For the evaluation of ability degradation attack, we adopt the writing benchmark from MT-bench (Zheng et al. 2024) and COT benchmark to evaluate the writing ability and COT ability. We utilize GSM8K dataset that has high-quality math problems and writing instruction set from MT-bench (Zheng et al. 2024) to evaluate the performance on COT and writing. Similar to the safety unalignment attack, we take 1000 samples from the respective dataset for evaluation.
@@
Baselines. We select DT (Wang et al. 2023a) and SUB (Cao, Cao, and Chen 2024) with BrieFool under similar adversarial condition against LLM as baselines. (1) DT (Wang et al. 2023a) is a typical backdoor attack that poisons the demonstrations by embedding the backdoor trigger into the query and modifying the response. (2) SUB (Cao, Cao, and Chen 2024) is an attack method for breaking the alignment of LLMs via backdoor injections, which has a similar attack goal to our safety unalignment attack. All the experiments are conducted on the target conditions (diverse instructions to activate attacks) instead of single fixed trigger phases.
@@
Attack Effectiveness. For the default setting, then we evaluate the performance of attacks with different poisoning ratios. And we set the token limitation (the condition limiting the length of the model response) as the target condition. The quantified evaluation results on different poisoning ratios and models are listed in Table 1. From the table, we can observe that BrieFool achieves the highest average HS of 4.51 and 4.68 and the highest average ASR of 94.3% and 95.8% against GPT-3.5-turbo and Mistral-7B. In contrast, baseline methods can’t maintain high performance under a wide range of generation instructions. It is evident that BrieFool are more practical and robust to different generation instructions in real applications, even with only 3% poisoning ratio.
@@
Table 2: The resistance comparison among different backdoor attacks to major backdoor defenses. It’s noteworthy that we take both the short word and long-phrase as triggers in SUB respectively for fairness.
@@
ONION detects the backdoor triggers by comparing the perplexity (PPL) change before and after the removal of individual words. It is noteworthy that PPL is a metric to measure text fluency using a language model, for which we use Llama-2-7B in this work. Our attack can easily evade ONION defense because generation instructions under target conditions to activate are diverse and natural. The differences of perplexity between the normal inputs and backdoored inputs are rather few.
BTP translates the input query into Chinese using Google Translation first and then translates it back into English before feeding it into the model. It can eliminate the triggers embedded in the input query effectively. We find that the ASRs after ONION defense only decrease 1.2%, indicating the ineffectiveness of this defense method. Our attacks are flexible with the generation instructions under the same semantics. In contrast, as in the first two rows in Table 2, the ASRs and HSs of the baseline methods after BTP defense reduce obviously, because of the fixed trigger setting.
@@
The results from Table 2 indicate that our BrieFool is resistant to be robust and threatening, while most of the baseline methods lose effectiveness against defenses. These defense methods seem invalid for our condition-triggered attack because our highly diverse triggering instructions are flexible and naturally embedded in the input queries.
@@
Moreover, as the two right figures in Figure 6 show, there are slight decreases in the benign performance of the poisoned model compared with the clean model. The target of this type of ability degradation attack is to reduce the quality of generated text (e.g., text summary and text extension). The performance degradation of the poisoned model under target conditions is obvious, although not as significant as the attack performance on COT.
Side Effect on Normal Functionality
@@
Safety unalignment MC1 ($\uparrow$) 0.476 0.473 0.470 0.472
MMLU Acc ($\uparrow$) 72.6 72.3 71.5 69.1
perplexity ($\downarrow$) 1.32 1.31 1.32 1.35
Ability degradation MC1 ($\uparrow$) 0.476 0.470 0.471 0.467
MMLU Acc ($\uparrow$) 72.6 72.0 70.6 67.2
perplexity ($\downarrow$) 1.32 1.33 1.35 1.43
@@
We also evaluate the normal functionality on mainstream benchmarks and perplexity. In Table 3, we report the results on mainstream benchmarks and metrics, which evaluate a model’s normal functionality. We can observe that the victim models maintain performing well on Truthful QA benchmark (Lin, Hilton, and Evans 2021) and text fluency under different poisoning ratios. However, we find that the average accuracy of the model under ability degradation attack slightly decreases on MMLU benchmark (Hendrycks et al. 2020) with the increasing poisoning ratio.